OpenStudy (anonymous):

I have the answer to this problem but I don't know how to get the answer. Please help. 36m^(4)n^(3)/(24m^(2)n^(5)) write in lowest terms. Answer: 3m^2/2n^2

OpenStudy (anonymous):

\[\frac{36m^4n^3}{24m^2n^5}\] first note \(\frac{36}{24}=\frac{3}{2}\) in lowest terms

OpenStudy (anonymous):

then \(\frac{m^4}{m^2}=m^{4-2}=m^2\) essentially you are cancelling \(m^2\) from top and bottom

OpenStudy (anonymous):

and finally \(\frac{n^3}{n^5}=\frac{1}{n^{5-3}}=\frac{1}{n^2}\) again this is cancelling, this time cancelling \(n^3\) tp and bottom

OpenStudy (anonymous):

put it all together gives \[\frac{36m^4n^3}{24m^2n^5}=\frac{3m^{4-2}}{2n^{5-3}}=\frac{3m^2}{2n^2}\]

OpenStudy (anonymous):

Oh thank you so much. I'll try these with my other problems now. But Theres a hareder one. If I post that one ould you walk me through that as well. I'll try the other stuff on my own. This helped alot.

OpenStudy (anonymous):

@satellite73